One acetic acid solution is 70 percent water, and another is 30 percent water. How many liters of each solution should be mixed to produce 20 liters of a solution that is 40 percent water?
Answer by stanbon(75887)

One acetic acid solution is 70 percent water, and another is 30 percent water. How many liters of each solution should be mixed to produce 20 liters of a solution that is 40 percent water?
--------------
Equation::
water + water = water
0.70x + 0.30(20-x) = 0.40*20
-----
70x + 30*20 - 30x = 40*20
----
40x = 30*20
---
x = 15 liters (amt. of 70% solution needed)
20-x = 5 liters (amt. of 30% solution needed)
------------
Cheers,
Stan H.
-------------